Considered a uniformly charged ring of radius R find the point on the axis where the electric field is maximum

Grade:12

1 Answers

Eshan
5 years ago
Dear student,

Electric field due to a charged ring at a distancexfrom it is given as

E=\dfrac{kQx}{(x^2+R^2)^{3/2}}
The derivative of E wrt x is zero at the point where the electric field is maximum.

Hence\dfrac{dE}{dx}=kQ(\dfrac{(x^2+R^2)^{3/2}.(1)-x\dfrac{3}{2}(x^2+R^2)^{1/2}.(2x)}{x^2+R^2})=0
\implies x=\dfrac{R}{\sqrt{2}}